9.Let A, B, C be the points (2,3)(3,−2) and (−1,4)
. The length of AB yields_____
a.square root of 26
b.square root of 27
c.square root of 23
d.square root of 35

10.____ is defined as the ratio of the vertical distance the line rises or falls to the horizontal distance.
a.Gradient of a line
b.Gradient of a helix
c.Gradient of a circle
d.None of the options
1
Expert's answer
2019-04-02T11:34:38-0400

9) "|AB|=\sqrt{(2-3)^2+(3+2)^2}=\sqrt{26}"

Answer a.square root of 26

10). Answer a.Gradient of a line